OpenStudy (anonymous):

1.Why do we need to cube root if the question want "mean separation"of particle or atom? 2. A spherical drop of oil of diameter of 0.90mm spreads out on the surface of some water to form a circular patch of diameter 18 cm. Assuming that the oil patch is one oil molecule thick, estimate the length (or mean diameter) of an oil molecule.

OpenStudy (unklerhaukus):

cube root might have something to do with volume

OpenStudy (unklerhaukus):

\[V_{sphere}=\frac 43\pi r^3\qquad\qquad A_{circle}=\pi r^2\qquad r=\frac d2\]